OpenStudy (anonymous):

helpppp. Two U.S. states are not part of the continental United States. What percent of the fifty U.S. states are included in the continental United States?

OpenStudy (anonymous):

96%

OpenStudy (anonymous):

48/50 states are included so 96%

OpenStudy (anonymous):

Total number of states = 50 Number of states not included = 2 Number of states included = 50 - 2 = 48 So, % of states included = 48/50*100 = 96%